OpenStudy (anonymous):

how far in miles would a school bus go in 4 hours, 25 minutes, at 110 kn/h? select the correct units also

OpenStudy (anonymous):

so how do you do that in a calualtor?

OpenStudy (anonymous):

You want to represent the hours as fractions of an hour: 4hr 25min = 265/60 hours d=vt=((265/60) hr)(110km/hr)

OpenStudy (anonymous):

In a calculator, 265*110/60

OpenStudy (anonymous):

\[485\frac{5}{6}km\]

OpenStudy (anonymous):

its not km or kg whats the correct units

OpenStudy (anonymous):

wouldnt it be miles since thats what they are looking for?

OpenStudy (anonymous):

oh, in miles, right lol. glanced over that

OpenStudy (anonymous):

Start by converting 4 hours, 25 minutes to hours, then determine how many kilometers the bus will travel using distance = (rate)(time).

OpenStudy (anonymous):

how you do it like that?

OpenStudy (anonymous):

\[485\frac{5}{6}km(\frac{0.621371192 mi}{1km}) \approx 301.883mi\]

OpenStudy (anonymous):

I showed you how to do it like that above, but i forgot to convert to miles afterwards

OpenStudy (anonymous):

404.1666667?

OpenStudy (anonymous):

what is that?

OpenStudy (anonymous):

485(5/6)

OpenStudy (anonymous):

\[485\frac{5}{6}\approx 485.833\]

OpenStudy (anonymous):

You don't multiply by the 5/6, the 5/6 is part of a mixed fraction

OpenStudy (anonymous):

\[485\frac{5}{6}=\frac{2915}{6}\]

OpenStudy (anonymous):

so 485.833 miles

OpenStudy (anonymous):

No, 485.833 km

OpenStudy (anonymous):

convert to miles, i showed that calculation above